OpenStudy (rissyroo118):

WILL MEDAL ! 2005 |2006| 2007 |2008 Plant 1 20 4 25 10 Plant 2 25 5 5 48 Plant 3 20 8 20 32 Plant 4 43 10 10 47 Plant 5 40 11 15 48 During which year were the mean and median the same value? A)2005 B)2006 C)2007 D)2008

OpenStudy (anonymous):

Well, to find the mean, you add up all the numbers and divide by the amount of numbers. And, to find the median you put the numbers from least to greatest and find the middle. (If there are 2 middles, you add them together and divide by 2) Im not going to do it for you, but i will help you find it.

OpenStudy (anonymous):

In 2005, the numbers from least to greatest are: 20, 20, 25, 40, 43

OpenStudy (rissyroo118):

so for the mean i got 446 so divide by how many numbers there are that would be 20 so 446/20 =22.3

OpenStudy (anonymous):

Actually, were working only on 2005 right now, so you would add 20 + 20 + 25 + 40 + 43 and divide it by 5. Try again :)

OpenStudy (rissyroo118):

128/5=25.6

OpenStudy (anonymous):

Correct! Great job! So, your mean in 2005 is 25.6 Now, we have to check if 25.5 is also the median.

OpenStudy (anonymous):

25.6*

OpenStudy (rissyroo118):

than 38/5=7.6 would be for 2006 than 75/5=15 for 2007 than 185/5=37 for 2008

OpenStudy (rissyroo118):

so now i should do the median for each year right?

OpenStudy (anonymous):

Excellent job rissyroo! Now, lets work on the medians :) Lets start by ordering them all from least to greatest.

OpenStudy (anonymous):

I did 2005 to get you started :) 2005: 20, 20, 25, 40, 43 2006: 2007: 2008:

OpenStudy (rissyroo118):

thank you :) so 2005: 20, 20, 25, 40 ,43 2006: 4, 5, 8, 10 ,11 2007: 5, 10 , 15 ,20 , 25 2008: 10 , 32 , 47 , 48 , 48

OpenStudy (anonymous):

Yep! Now, you can use your finger if you like (I do!) to find the number in the middle! (If there are 2 numbers in the middle, add them together and divide by 2)

OpenStudy (anonymous):

For example, the median of 2005 would be 25 because its in the middle :)

OpenStudy (rissyroo118):

8 & 15 i think right because we cancel out 2005 & 2008? find the number from 2006 & 2007. 8+15=23 23/2=11.5

OpenStudy (anonymous):

actually, there all 5's and 5 is an odd number, so there would be a middle.. the medians are below if this helps: Median: 25 = 2005 8 = 2006 15 = 2007 47 = 2008

OpenStudy (anonymous):

if you had 6 numbers, you would add the 2 in the middle to find your middle, but these are all in 5's so there is a middle for each data set.

OpenStudy (anonymous):

So, im placing the means and medians below. Tell me which year shares the same number :) Mean: 25.6 = 2005 7.6 = 2006 15 = 2007 37 = 2008 Median: 25 = 2005 8 = 2006 15 = 2007 47 = 2008

OpenStudy (rissyroo118):

i understand much better now since you broke it down and explained too me. thank you! the answer would be 2007

OpenStudy (anonymous):

Exactly! Very good! No problem rissyroo118, if you have any more problems just tag me and ill try to help the best i can :) You did a very good job! i hope this helps with your review/homework/test/etc.!